# Secure Data Reconstruction: A Direct Data-Driven Approach

### v1

@@ III Entry-Attacked Scenario
Lemma 1 is related to the fundamental lemma [16]. However, it does not require an input/output partitioning of the variables nor controllability of the system $\mathscr{B}$. Instead of a condition on an input component of the trajectory (persistentcy of excitation of an input), it imposes a condition on the data $w_{d}$ (generalized persistency of excitation). Finally, the conditions of Lemma 1 are necessary and sufficient while the ones of [16] are only sufficient [18].
@@ III-B A brute force algorithm
It is true that when fewer than $k$ entries are adversarial, we may end up discarding more data than necessary. To mitigate this, we can start with a smaller value $\bar{k}<k$. If no solution is found for a given $\bar{k}$, it indicates that either $\bar{k}$ is too small (i.e., we are not removing enough entries) or the data recovery problem itself is infeasible (see Theorem 1 later). In this case, we iteratively increase $\bar{k}$ until Algorithm 1 finds a solution, or until $\bar{k}=k$. If the algorithm still fails to find a solution at the upper bound $k$, we conclude that recovering the true trajectory is unlikely.
@@ III-C Performance analysis
Clearly, if $\operatorname*{\mathcal{I}}^{(i)}=[qL]/\mathcal{C}$, meaning all compromised entries are exactly removed, we can solve (5) and Algorithm 1 yields the true trajectory $\tilde{w}=\bar{w}$ [13, Proposition 9]. However, a natural question arises: if (5) is solvable, does this necessarily imply that $\operatorname*{\mathcal{I}}^{(i)}=[qL]/\mathcal{C}$ and $\tilde{w}=\bar{w}$? More generally, what conditions guarantee that the output of Algorithm 1 always coincides with the real trajectory? This subsection is dedicated to addressing these questions.
@@
Suppose that no more than $k$ entries are malicious, (1) is satisfied, and Condition 1 holds. If Algorithm 1 finds a solution $\tilde{w}$, then it follows that $\tilde{w}=\bar{w}$. That is, the real trajectory is recovered.
@@ IV-B Performance analysis
Suppose that no more than $k$ channels are malicious, (1) is satisfied, and Condition 2 holds. If Algorithm 2 finds a solution $\tilde{w}$, then it follows that $\tilde{w}=\bar{w}$. That is, the real trajectory is recovered.
@@
Suppose that no more than $k$ channels are malicious, (1) is satisfied, but Condition 2 does not holds. Even if Algorithm 2 yields a solution $\tilde{w}$, it is possible that $\tilde{w}\neq\bar{w}$.
@@
Based on the above results, Condition 2 is crucial to well-define the problem. That is, in order to ensure the recovery of the true trajectory, removing the rows corresponding to any $2k$ channels must not compromise the rank of $\operatorname*{\mathscr{H}}$. Otherwise, the problem becomes infeasible. Moreover, for the same reason discussed in Remark 2, it is essential in working against the worst-case attacks.
@@ V-A Approximation solutions
Many algorithms are available in the literature to solve the $l_{1}$ norm minimization problem (22) and the group LASSO problem (33), including coordinate descent, exact path-following, and block coordinate gradient descent methods [31]. The computational complexities of these algorithms are typically polynomial with respective to the number of channels, the size of the Hankel matrix, and the length of the trajectory. In contrast, brute-force algorithms show exponential complexity. such as the coordinate descent, exact path-following, and block coordinate gradient descent methods. Therefore, (22) and (33) significantly reduce the computation costs.

### v2

@@ III Entry-Attacked Scenario
Lemma 1 is related to the fundamental lemma [16]. However, it does not require an input/output partitioning of the variables nor controllability of the system $\mathscr{B}$. Instead of a condition on an input component of the trajectory (persistency of excitation of an input), it imposes a condition on the data $w_{d}$ (generalized persistency of excitation). Finally, the conditions of Lemma 1 are necessary and sufficient while the ones of [16] are only sufficient [18].
@@ III-B A brute force algorithm
It is true that when fewer than $k$ entries are adversarial, we may end up discarding more data than necessary. To mitigate this, we can start with a smaller value $\bar{k}<k$. If no solution is found for a given $\bar{k}$, it indicates that $\bar{k}$ is too small, that is, we are not removing enough entries. In this case, we iteratively increase $\bar{k}$ until Algorithm 1 finds a solution.
@@ III-C Performance analysis
Algorithm 1 searches through all possible sets $\mathcal{I}^{(i)}$ such that $|\mathcal{I}^{(i)}|=qL-k$. Among these, there must exist a set $\mathcal{I}^{(i)}$ that contains only the indices of benign entries, that is, $\operatorname*{\mathcal{I}}^{(i)}\subset[qL]/\mathcal{C}$, as the total number of benign entries is no less than $qL-k$. In this case, all compromised entries are removed. Therefore, we can always solve (5) and Algorithm 1 produces the true trajectory $\tilde{w}=\bar{w}$ [13, Proposition 9]. However, a natural question arises: when (5) yields a solution, does this necessarily imply that we have excluded all malicious entries and $\tilde{w}=\bar{w}$? More generally, what conditions guarantee that the output of Algorithm 1 always coincides with the real trajectory? This subsection is dedicated to addressing these questions.
@@
Suppose that no more than $k$ entries are malicious, (1) is satisfied, and Condition 1 holds. When Algorithm 1 finds a solution $\tilde{w}$, it follows that $\tilde{w}=\bar{w}$. That is, the real trajectory is recovered.
@@ IV-B Performance analysis
Suppose that no more than $k$ channels are malicious, (1) is satisfied, and Condition 2 holds. When Algorithm 2 finds a solution $\tilde{w}$, it follows that $\tilde{w}=\bar{w}$. That is, the real trajectory is recovered.
@@
Suppose that no more than $k$ channels are malicious, (1) is satisfied, but Condition 2 does not hold. Even if Algorithm 2 yields a solution $\tilde{w}$, it is possible that $\tilde{w}\neq\bar{w}$.
@@
Based on the above results, Condition 2 is crucial to well-define the problem. That is, in order to ensure the recovery of the true trajectory, removing the rows corresponding to any $2k$ channels must not compromise the rank of $\operatorname*{\mathscr{H}}$. Otherwise, the problem becomes infeasible. Moreover, for the same reason discussed in Remark 2, this condition is essential in working against the worst-case attacks.
@@ V-A Approximation solutions
Many algorithms are available in the literature to solve the $l_{1}$ norm minimization problem (22) and the group LASSO problem (33), including coordinate descent, exact path-following, and block coordinate gradient descent methods [31]. The computational complexities of these algorithms are typically polynomial with respective to the number of channels, the size of the Hankel matrix, and the length of the trajectory. In contrast, brute-force algorithms show exponential complexity. Therefore, (22) and (33) significantly reduce the computation costs.